OpenStudy (hang254):

Find the inverse of the function. f(x) = x^3 + 5

OpenStudy (hang254):

@ganeshie8 @cwrw238

OpenStudy (jdoe0001):

.... hmmm do you know how to find inverse functions?

OpenStudy (hang254):

i think it is \[f^-1(x) = \sqrt[3]{x-5}\]

OpenStudy (hang254):

to a certain extent

OpenStudy (jdoe0001):

right, however ... one sec

OpenStudy (jdoe0001):

notice the graph of the resulting "expression" do you think it will pass the "vertical line test" needed to have it as a function? so, you can simplify the expression, it just won't give a "function as defined"

OpenStudy (jdoe0001):

so \(\bf f(x) = x^3 + 5 \) has no inverse function

OpenStudy (hang254):

oh, so i was incorrect.

OpenStudy (hang254):

thanks! @jdoe0001

OpenStudy (jdoe0001):

yw

OpenStudy (agent0smith):

"Find the inverse" to me means find the inverse, whether it's a function or not @jdoe0001

OpenStudy (agent0smith):

And x^3+5 is one to one and thus has an inverse which is a function...

OpenStudy (agent0smith):

https://www.google.com/search?q=(x-5)%5E(1%2F3)&oq=(x-5)%5E(1%2F3)&aqs=chrome..69i57j0l5.6401j0j7&sourceid=chrome&espv=210&es_sm=93&ie=UTF-8 inverse which is a function. Idk what your graph is @jdoe0001 but it's not correct.

OpenStudy (agent0smith):

@hang254 you were correct with \[\Large f^{-1}(x) = \sqrt[3]{x-5}\]

OpenStudy (agent0smith):

If a function passes the horizontal line test, as x^3+5 does, then it has an inverse which is a function.